archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    Details The VPS GEO infrastructure of Dallas TX is being tested and waiting for installation of the routers which must be there before the end of the month Date Saturday 05 May 2012 03 16AM Reason for closing Done Comments

    Original URL path: http://status.ovh.es/?do=details&id=1971 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    CLOSED Percent Complete Details The VPS GEO infrastructure of Madrid is being tested The expedition is planned for Thursday Date Saturday 05 May 2012 03 15AM Reason for closing Done Comments 2 Comment by OVH Thursday 27 October 2011 10 34AM The infrastructure is being deployed and we should have the service UP around Friday night Comment by OVH Saturday 29 October 2011 14 15PM The infrastructure is UP We

    Original URL path: http://status.ovh.es/?do=details&id=1968 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Status CLOSED Percent Complete Details The VPS GEO equipments of Milan are being tested The expedition is planned for Thursday Date Saturday 05 May 2012 03 15AM Reason for closing Done Comments 2 Comment by OVH Thursday 27 October 2011 10 32AM The infrastructure is being installed at Milan Tomorrow the service should be UP Comment by OVH Saturday 29 October 2011 14 13PM The infrastructure is UP We are

    Original URL path: http://status.ovh.es/?do=details&id=1967 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details The GEO VPS infrastructure of Frankfurt are being tested All must go to Frankfurt Wednesday Date Saturday 05 May 2012 03 15AM Reason for closing Done Comments 1 Comment by OVH Wednesday 26 October 2011 07 37AM

    Original URL path: http://status.ovh.es/?do=details&id=1966 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    to Project VPS Task Type Modernization Category Infrastructure Status CLOSED Percent Complete Details The VPS GEO infrastructure in Paris has been deployed We still have to change the switchs and re test all of them Date Saturday 05 May 2012

    Original URL path: http://status.ovh.es/?do=details&id=1965 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Reason for closing Done Additional comments about closing The problem has been fixed Everything is back to normal stage For information The vcenter VPS contains many hosts filers vms its rebooting takes between 20 to 30mn so that the system lists all available resources Besides There had been a problem on HA cluster In fact any request concerning this cluster caused a time out at the vcenter level So we

    Original URL path: http://status.ovh.es/?do=details&id=1936 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    to Project VPS Task Type Maintenance Category Infrastructure Status CLOSED Percent Complete Details We will update NX OS pf the VPS infrastructure Date Saturday 08 October 2011 09 04AM Reason

    Original URL path: http://status.ovh.es/?do=details&id=1905 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    the A side vps 10 n5 vers vps 20 n5 Comment by OVH Friday 07 October 2011 16 25PM The A side is cut on the vps 10 n5 We disconnect on the vps 20 n5 Comment by OVH Friday 07 October 2011 16 26PM A host is not correctly connected on the B side Comment by OVH Friday 07 October 2011 16 28PM The A side is connected on the vps 20 n5 We attack the B side ie we upgrade FEX of vps 11 n5 on vps 21 n5 We cut the B port on host Comment by OVH Friday 07 October 2011 16 46PM We have forgotten to reconfigure the router Fixed Comment by OVH Friday 07 October 2011 16 47PM Everything is cut on the B side Every thing work on the A side Comment by OVH Friday 07 October 2011 16 48PM We have finalised the B side it will take roughly 2h Comment by OVH Friday 07 October 2011 16 49PM We switch on the B side Comment by OVH Friday 07 October 2011 16 53PM The B side is up We have finalized the setting of 4x10G per rack We switch the routing configuration on the rbx s10 6k So we will cut traffic on the s10 It s going to work on the s9 Then we change the configuration Comment by OVH Friday 07 October 2011 16 55PM The B side is up We have finalized the setting of 4x10G per rack We switch the routing configuration on the rbx s10 6k So we will cut traffic on the s10 It s going to work on the s9 Then we change the configuration Comment by OVH Friday 07 October 2011 16 56PM The traffic is cut on the s10 Everything is routed

    Original URL path: http://status.ovh.es/?do=details&id=1902 (2015-10-15)
    Open archived version from archive



  •